# CoSteer: Collaborative Decoding-Time Personalization via Local Delta Steering

### v2

@@ Abstract
Personalized text generation has become crucial for adapting language models to diverse and evolving users’ personal context across cultural, temporal, and contextual dimensions. While existing methods often rely on centralized fine-tuning or static preference alignment, they struggle to achieve real-time adaptation under resource constraints inherent to personal devices. This limitation creates a dilemma: large cloud-based models lack access to localized user-specific information, while small on-device models cannot match the generation quality of their cloud counterparts. To address this dichotomy, we present CoSteer, a novel collaborative framework that enables decoding-time personalization through localized delta steering. Our key insight lies in leveraging the logits difference between personal context-aware and -agnostic outputs from local small models as steering signals for cloud-based LLMs. Specifically, we formulate token-level optimization as an online learning problem, where local delta vectors dynamically adjust the remote LLM’s logits within the on-device environment. This approach preserves privacy by transmitting only the final steered tokens rather than raw data or intermediate vectors, while maintaining cloud-based LLMs’ general capabilities without fine-tuning. Through comprehensive experiments on various personalized generation tasks, we demonstrate that CoSteer effectively assists LLMs in generating personalized content by leveraging locally stored user profiles and histories, ensuring privacy preservation through on-device data processing while maintaining acceptable computational overhead.
11footnotetext: Equal Contribution.11footnotetext: Corresponding authors.
@@ 1 Introduction
The rapid development of large language models (LLMs) has significantly enhanced natural language processing, enabling these models to understand context and generate coherent text effectively (Zhao et al., 2023; Gu et al., 2025; Huang et al., 2025). This progress has sparked interest in personalization, where AI systems move beyond generic content to tailor interactions based on individual user profiles. Personalized generation refers to creating content customized through analysis of user-specific attributes like linguistic patterns, interaction histories, and contextual preferences (Xu et al., 2025; Liu et al., 2025a). This approach enables user-specific outputs that maintain contextual relevance, with applications in personalized recommender systems (Lyu et al., 2023; Zhang et al., 2024a; Yu et al., 2025), adaptive dialogue agents (Wu et al., 2025), and customized content creation platforms (Mysore et al., 2024).
Existing personalized generation approaches primarily fall into two paradigms. The first involves training-based methods, which leverage user data to tailor models using two main strategies. Parameter-efficient architectures, such as PLoRA (Zhang et al., 2025a), employ a shared LoRA module across users to strike a balance between personalization and resource efficiency. In contrast, individualized adapters, as seen in UserAdapter (Zhong et al., 2021), optimize for per-user customization, albeit at a higher computational cost. Additionally, multi-objective reinforcement learning frameworks (Zhou et al., 2024; Wu et al., 2023) redefine personalization as the alignment of generic model capabilities with user-specific patterns, achieved through reward shaping. The second paradigm comprises tuning-free methods, which avoid model updates by adapting to user context during inference. This approach is further divided into two techniques. Prompt engineering implicitly encodes user profiles through curated prompts. For example, Cue-CoT (Wang et al., 2023a) integrates user-specific reasoning into system prompts, while PAG (Richardson et al., 2023) retrieves historical interactions to construct context-aware prompts. On the other hand, inference-time optimization techniques, exemplified by PAD (Chen et al., 2025) and Drift (Kim et al., 2025), actively adjust token generation probabilities by manipulating logit differences. Unlike prompt-based methods that inject semantic context, these techniques constrain output distributions to better align with user preferences.
@@
However, both training-based and tuning-free methods face critical dilemmas in balancing privacy and quality when deployed on resource-constrained personal devices. Training-based methods require computational resources that may not be available locally, and cloud-based training risks exposing user data. Meanwhile, tuning free methods that transmit raw personal context to cloud-based LLMs for context-aware adaptation also risk exposing sensitive information, whereas relying solely on local SLMs leads to degraded output quality. Therefore, there is a pressing need to design a new framework that balances quality and privacy by combining the general capabilities of remote LLMs with the benefits of local information. Achieving this balance on personal devices faces two major challenges: 1)User information changes in real-time, necessitating methods for real-time adjustments to maintain personalization effectively. 2)To avoid the risk of user information leakage, any privacy-sensitive operations must be conducted locally.
Figure 1: Schematic illustration of CoSteer framework. (a) Task scenario: A user poses a question potentially requiring access to local personal context (e.g., user profile, interaction history). (b) Limitations of small locally-deployed language models: Direct inference with constrained model capability leads to suboptimal generation quality. (c) Challenges of cloud-based LLMs: Despite strong generalization, once LLMs are constrained from accessing local personal context, they result in misaligned or contextually disconnected outputs. (d) CoSteer: Optimizes LLM predictions through local delta steering, effectively balancing the LLM’s broad knowledge with user-specific information.
@@
To address these fundamental challenges, we propose CoSteer, a collaborative framework enabling real-time personalized generation via localized delta steering during decoding. Our core innovation resides in utilizing the logits difference, i.e., delta, between personal context-aware and -agnostic outputs from on-device small language models (SLMs) as dynamic steering signals to guide cloud-based LLM distributions. Specifically, the SLM generates two contrasting predictions: 1) personalized outputs incorporating privacy-sensitive personal context, and 2) generic outputs using users’ queries only. The resultant logit differentials serve as indicators of personalization directions.
By formulating token-level adaptation as an online learning process, CoSteer establishes an iterative refinement mechanism that enables edge devices to locally optimize cloud-based LLM predictions, eliminating the need to transmit raw personal context or intermediate representations. Experimental evaluations demonstrate that CoSteer achieves superior performance compared to two conventional baselines: 1) cloud-based LLMs processing queries without personalization context, and 2) local SLMs conducting personalization context-aware inference independently. Remarkably, the framework attains comparable performance to the near-upper-bound baseline of cloud-based LLMs operating with full personal context access, while maintaining strict data isolation.
To sum up, our contributions are threefold:
  • Pioneering Focus on Personalized Generation in Resource-Constrained Environments: We are among the first to address the issue of personalized generation in edge environments with limited resources, a problem of urgent research significance in today’s context.
  • Introduction of the CoSteer Framework: We propose the CoSteer framework, which optimizes LLM’s predictions using local delta signals. This approach effectively combines the general capabilities of LLMs with the privacy benefits of on-device SLMs.
  • Comprehensive and Realistic Validation: We provide substantial experimental evidence demonstrating the effectiveness of our proposed method. Our evaluation further extends to address critical real-world challenges, confirming the framework’s practical value.
@@ 2 Related work
Training-time personalization
@@
Current mainstream approaches to personalization predominantly focus on training-time adaptation, leveraging user-specific data to tailor models through parameter-efficient fine-tuning (PEFT) or reinforcement learning-based alignment. A subset of work explores one-PEFT-for-all-users strategies (Zhang et al., 2025a; Woźniak et al., 2024; Zhu et al., 2024; Kong et al., 2025), where a shared set of lightweight parameters (e.g., adapters or conditional batch normalization modules) is optimized to generalize across diverse user preferences. Conversely, the one-PEFT-per-user paradigm prioritizes individualized tuning through isolated parameter sets, enhancing personalization while preserving privacy by avoiding cross-user data leakage (Zhong et al., 2021; Peng et al., 2024; Tan et al., 2025; 2024). Beyond PEFT, RL-based approaches have gained traction for aligning models with user preferences through reward-driven optimization. Recent work has further formalized personalization as a multi-objective reinforcement learning (MORL) problem: methods like MORLHF (Wu et al., 2023) and MODPO (Zhou et al., 2024) train distinct reward models for different objectives and merge them during optimization, while alternative frameworks such as Personalized Soups (Jang et al., 2023), Reward Soups (Ramé et al., 2023), MOD (Shi et al., 2024), and PAD (Chen et al., 2025) dynamically combine policies from multiple trained models during the decoding phase. This series of work requires significant computational resources and relies on static datasets, which cannot meet the real-time personalization needs of users.
Inference-time adaptation
@@
Methods like Linear Alignment (Gao et al., 2024)and Contrastive Decoding (Li et al., 2023) pioneer logit steering mechanisms that dynamically adjust token distributions during generation. By computing differential signals between personalized and generic outputs on edge devices, these approaches eliminate dependency on retraining while preserving privacy through localized computation. CoS (He et al., 2025) follows this approach by using an adaptable parameter $\lambda$ to scale the impact of personal information, thereby achieving controllable personalization. Amulet (Zhang et al., 2025b) uses online learning algorithms to iteratively optimize the logits distribution, precisely aligning with the user’s preferences. However, this series of methods requires explicitly transmitting user information to the LLM, which poses a risk of privacy leakage (Yan et al., 2025) and hinders the practical deployment of these methods.
Collaborative generation
@@
The computational constraints of edge devices have driven innovative paradigms in collaborative generation between cloud-based LLMs and local small language models. Existing collaborative generation approaches predominantly focus on two technical routes: (1) assistive alignment, where local SLMs are trained to augment the LLM’s user preference alignment during inference, and (2) reward-guided decoding, where lightweight reward models provide preference signals through reward-guided decoding. For instance, Aligner (Ji et al., 2024) leverages natural language feedback generated by the SLM to directly inject user-specific linguistic patterns into the LLM’s decoding process, while Expo (Zheng et al., 2025) achieves preference alignment through linear interpolation of layer-wise weights between LLMs and SLMs. Recent advancements further demonstrate that training specialized small reward models can effectively steer LLM outputs toward personalized objectives (Snell et al., 2024; Liu et al., 2025b). Among these, Proxy-tuning (Liu et al., 2024) and Cogensis (Zhang et al., 2024b) are closest to our design. Proxy-tuning modifies the LLM’s logits distribution by contrasting pre-trained and fine-tuned SLM outputs, whereas Cogensis employs a learned fusion network to combine logits from both models. In contrast to these existing methods, our proposed CoSteer framework establishes a completely tuning-free collaboration mechanism, eliminating the need for training.
@@ 3.1.1 Task formulation
Our research addresses the challenge of personalized text generation in resource-constrained local environments. This scenario involves a dual-model architecture where a cloud-based LLM processes only the raw user query, and a locally deployed SLM has access to privacy-sensitive personal context, including user preferences, profiles, and interaction histories.
The primary objective is to optimize generation quality by synergizing the LLM’s general linguistic capabilities with localized personalization while maintaining strict privacy preservation. Unlike conventional approaches that frame decoding as a continuous Markov Decision Process (MDP), we reformulate per-token generation as an independent online optimization task. In this paradigm, the LLM’s logit distribution during decoding serves as the policy $\pi$, enabling us to formalize the optimization objective for per-token generation as:
$\pi^{*}(a)=\underset{\pi\in\Pi}{\arg\max}\mathbb{E}_{a\sim\pi\left(\cdot\mid p,s\right)}r\left(a\mid p,s\right)$ (1)
where $p$ denotes the initial prompt, $s$ represents the sequence that has already been generated, $a\in\mathcal{A}$ is a potential token in the vocaburary space $\mathcal{A}$, and $r$ is the latent reward function that reflects the current user’s personal context.
@@ 3.1.2 Online learning
User personalization tasks are characterized by individuality, diversity, and changing over time. Traditional offline learning methods rely on static datasets, which struggle to capture dynamic personalized information in real-time. Online learning, however, enables the model to adapt instantly using the latest personalized context.
Inspired by the recent work Amulet (Zhang et al., 2025b), our study employs the widely studied Follow-The-Regularized-Leader (FTRL) (McMahan, 2011) as the online learning algorithm. The FTRL is the FTL algorithm with a regularizer term added, which significantly improves both algorithmic stability and convergence properties. Typically, the policy optimization process of FTRL at iteration $t$ can be mathematically represented as:
$\pi_{t}(a)=\arg\max_{\pi\in\Pi}\left[\sum_{i=0}^{t-1}\mathcal{U}_{i}(\pi_{i}(a))-\frac{1}{\eta}\left(\phi(\pi)+\frac{\lambda}{2}\lVert\pi-\pi_{t-1}\rVert^{2}\right)\right]$ (2)
Given the practical constraint of inaccessible true user rewards, each iteration $t$ necessitates the employment of an approximate utility function $U$ to progressively estimate personalization signals for iterative policy refinement. The first term in Equation 2 implements the fictitious play mechanism to minimize regret between current and historical policies, while the second term represents the regularization and smoothness component with $\eta,\lambda$ as tunable hyperparameters.
@@ 3.2 Decoding-time personalization via local delta steering
From Equation 1, we understand that the distribution of logits is the strategy we need to optimize. Once we identify an appropriate utility function $U$ to characterize the relative quality of the current logits distribution, we can use Equation 2 to iteratively optimize it.
Series of recent works, including Contrastive Decoding (Li et al., 2023) and Linear Alignment (Gao et al., 2024), have demonstrated that the difference in logits before and after incorporating specific context into the model can serve as a direction for optimization (He et al., 2025; Zhang et al., 2025b). Thus from the online learning perspective, their utility function can be expressed as follows:
$u_{t}(a)=\alpha\left(\log\pi_{t}(a)-\log\pi_{\text{base }}(a)\right)$ (3)
However, this approach is not feasible in our task scenario, as it requires transmitting personal context to the cloud-based LLM, which poses a risk of privacy leakage. Inspired by this idea, we notice that both the small and LLMs share the same state space during decoding. Therefore, we can use delta, the difference in the SLM’s logits before and after incorporating personal context, to steer the LLM’s logits distribution. Formally, we define the utility function as follows:
@@
$u_{t}(a)=\underbrace{\alpha(\log\pi_{t}(a)-\log\pi_{base}(a))}_{\text{LLM Policy Contrast}}+\underbrace{\beta(\log\pi^{*}_{pers}(a)-\log\pi^{*}_{base}(a))}_{\text{SLM Delta Steering}}$ (4)
Here, we define $p_{base}$ as the user’s basic query, and $p_{pers}$ represents the privacy-sensitive personal context. We formalize three distinct generation policies:
  • Target Policy $\pi_{t}(a)$ that is being optimized at step $t$
  • LLM Base Policy with query only: $\pi_{base}(a)=P_{LLM}(a|p_{base},s)$
  • SLM Reference Policies:
    • $\pi^{*}_{base}(a)=P_{SLM}(a|p_{base},s)$: SLM baseline policy with query only
    • $\pi^{*}_{pers}(a)=P_{SLM}(a|p_{base},p_{pers},s)$: SLM policy with full personal context
where $s$ denotes the current token sequence, $P_{LLM}$$\pi$ denotes the LLM’s distribution, and $P_{SLM}$$\pi^{*}$ the SLM’s distribution. Hyperparameters $\alpha,\beta$ adjust the influence of each optimization component.
The utility function implements a dual-contrastive mechanism: The first term amplifies desirable deviations from the LLM’s base policy, while the second term aligns optimization with the personal context-aware SLM policy. Through iterative application, this formulation progressively accentuates the semantic differential induced by $p_{pers}$ relative to $p_{base}$, driving the target policy $\pi_{t}$ toward maximal utilization of local delta signals.
Additionally, to retain the LLM’s stronger general capabilities, we want to ensure that the final strategy does not deviate too far from the LLM’s initial strategy. Therefore, we incorporate a KL divergence term into Equation 4 to impose this constraint.
$\mathcal{U}_{t}(\pi)=u_{t}(\pi)-\lambda D_{\mathrm{KL}}\left(\pi\|\pi_{0}\right)$ (5)
where the initial policy $\pi_{0}=\pi_{base}$. The FTRL algorithm requires adding a regularizer to stabilize the algorithm. Here, we continue to use KL divergence as this term, leading to a FTRL-proximal-like iteration dynamic.
$\pi_{t}=\underset{\pi\in\Pi}{\arg\max}\left[\sum_{i=0}^{t-1}\mathcal{U}_{i}(\pi)-\frac{1}{\eta}D_{\mathrm{KL}}\left(\pi\|\pi_{t-1}\right)\right]$ (6)
By substituting Equation 4 into 5, and then 5 into 6, we can iteratively optimize the strategy. However, the current iterative computation incurs significant overhead, which is not feasible for our personalization scenarios that require rapid response. Therefore, we deduce the closed-form solution of Equation 6, as shown in Equation 7, to significantly reduce the additional computational cost. The derivation of the closed-form solution can be found in Appendix A.3.
3.3 CoSteer framework
After explaining how we use local delta signals to steer the target policy, let us formally introduce our CoSteer framework, as shown in Algorithm 1. During the inference of each token, the cloud-based LLM sends the current logits to the local environment. The local SLM simultaneously performs inference using prompts with and without personal context to obtain delta signals. Upon receiving the LLM’s logits from the cloud, the local environment applies Equation 7 for iterative optimization, samples the output token, and uploads it back to the LLM for inferring the next token. Through this method, CoSteer integrates the large language model’s general linguistic capabilities with localized personalization that maintains privacy preservation.

### v3

@@ Abstract
Personalization has become crucial for adapting models to the diverse and evolving needs of users across cultural, temporal, and contextual dimensions. While existing methods often rely on centralized fine-tuning or static preference alignment within a single model, they struggle to achieve both real-time and high-quality personalization under the resource and privacy constraints of personal devices. To address this challenge, we propose CoSteer, a collaborative framework that enables tuning-free, real-time personalization via decoding-time adaptation. By leveraging logit differences between context-aware and -agnostic local small models, CoSteer steers cloud-based large models, ensuring effective personalization while preserving the large model’s capabilities. Personalization is handled locally, with only final tokens sent to the cloud, maintaining both user context and system efficiency. Through extensive experiments across a wide range of tasks, we demonstrate that CoSteer generates high-quality personalized content, ensuring both effectiveness and computational efficiency. Our results highlight its robustness across models and environments, confirming its practical applicability in real-world scenarios.
Machine Learning, ICML
@@ 1 Introduction
The rapid development of large language models (LLMs) has significantly enhanced natural language processing, enabling these models to understand context and generate coherent text effectively (zhao2023llmsurvey). This progress has sparked interest in personalization, where AI systems move beyond generic content to tailor interactions based on individual user profiles. Personalized generation refers to creating content customized through analysis of user-specific attributes like linguistic patterns, interaction histories, and contextual preferences (Xu2025PersonalizedGI). This approach enables user-specific outputs that maintain contextual relevance, with applications in personalized recommender systems (lyu2023llmrec; zhang2024agentrec), adaptive dialogue agents (dialogagent), and customized content creation platforms (mysore2024pearlpersonalizinglargelanguage).
Existing personalized generation approaches primarily fall into two main paradigms. The first involves training-based methods, which use user data to tailor models, such as parameter-efficient architectures (plora) and individualized adapters (zhong-etal-2021-useradapter), as well as multi-objective reinforcement learning frameworks (modpo; morlhf) that align model capabilities with user-specific patterns. The second paradigm comprises tuning-free methods, which adapt to user context during inference without model updates, including prompt engineering (cuecot; pag) and inference-time optimization techniques (pad; Kim2025DriftDP) that adjust model output distribution based on specific user preference.
@@
However, both paradigms face critical challenges in balancing privacy and quality when deployed on resource-constrained personal devices. Training-based methods require significant computational resources, which may not be available locally, and cloud-based training poses risks of exposing user data. In contrast, tuning-free methods that transmit personal context to cloud-based LLMs for context-aware adaptation also risk data leakage, while relying solely on local compact models compromises output quality.
Therefore, there is a pressing need to design a new framework that synergizes remote LLM capabilities with local privacy. Achieving this balance on resource-constrained personal devices, however, requires overcoming two primary obstacles. First, user information evolves in real-time, necessitating dynamic adjustment methods to maintain personalization effectiveness without reliance on frequent, costly retraining. Second, a fundamental difficulty lies in effectively guiding the remote LLM to incorporate personalized patterns without ever transmitting the raw context to the cloud, demanding a mechanism that ensures precise personalization across strict privacy boundaries.
@@
To address these fundamental challenges, we propose CoSteer, a collaborative framework enabling real-time personalized generation via localized delta steering . Our core innovation resides in utilizing the logits difference between personal context-aware and -agnostic outputs from on-device small language models (SLMs) as dynamic steering signals to guide cloud-based LLM distributions. Specifically, the SLM generates two contrasting predictions: 1) personalized outputs incorporating privacy-sensitive personal context, and 2) generic outputs using users’ queries only. The resultant logit differentials serve as indicators of personalization directions. To effectively integrate these instantaneous steering signals into the sequential generation process without disrupting the LLM’s coherence, we mathematically formulate the decoding phase as an online learning problem. By treating token-level adaptation as an iterative optimization process, CoSteer establishes a refinement mechanism that enables edge devices to locally optimize cloud-based LLM predictions, eliminating the need to transmit raw personal context or intermediate representations.
Extensive evaluations across a diverse range of personalized tasks validate CoSteer’s versatility. Our results demonstrate that the framework consistently outperforms conventional baselines, achieving a remarkable “weak-to-strong” generalization where a tiny local model effectively steers a massive cloud LLM. Beyond performance gains, our comprehensive analysis confirms CoSteer’s practicality for real-world deployment: it facilitates seamless collaboration across heterogeneous model architectures, demonstrates strong robustness against noisy user contexts, and significantly reduces computational overhead compared to full-context cloud processing. Our contributions could be summarized as follows:
  • Personalization in Edge Environments: We tackle the challenge of deploying personalized generation on resource-constrained devices. Our approach addresses the inherent trade-off between privacy and quality, offering a solution that utilizes local context to enhance remote LLMs without exposing sensitive user data.
  • The CoSteer Framework via Online Learning: We propose CoSteer, a tuning-free framework that reformulates decoding-time adaptation as an online learning process. By utilizing local logit deltas to iteratively steer cloud predictions, it enables precise personalization while maintaining strict data isolation.
  • Validation of Practicality and Generalization: Extensive experiments demonstrate our capability, where a tiny local model can effectively steers a massive cloud LLM. We further confirm the framework’s real-world viability through its proven efficiency, robustness to noisy context, and cross-architecture compatibility.
@@ 2 Related work
Training-time Personalization.
@@
Current mainstream approaches to personalization predominantly focus on training-time adaptation, leveraging user-specific data to tailor models through parameter-efficient fine-tuning (PEFT) or reinforcement learning-based alignment. A subset of work explores one-PEFT-for-all-users strategies (plora; lmp; reclora; ilora), where a shared set of lightweight parameters (e.g., adapters or conditional batch normalization modules) is optimized to generalize across diverse user preferences. Conversely, the one-PEFT-per-user paradigm prioritizes individualized tuning through isolated parameter sets, enhancing personalization while preserving privacy by avoiding cross-user data leakage (zhong-etal-2021-useradapter; peng2024pocketllmenablingondevicefinetuning; tan2025democratizinglargelanguagemodels; tan2024personalizedpiecesefficientpersonalized). Beyond PEFT, RL-based approaches have gained traction for aligning models with user preferences through reward-driven optimization. Recent work has further formalized personalization as a multi-objective reinforcement learning (MORL) problem: methods like MORLHF (morlhf) and MODPO (modpo) train distinct reward models for different objectives and merge them during optimization, while alternative frameworks such as Personalized Soups (jang2023personalizedsoupspersonalizedlarge), Reward Soups (ramé2023rewardedsoupsparetooptimalalignment), MOD (mod), and PAD (pad) dynamically combine policies from multiple trained models during the decoding phase. This series of work requires significant computational resources and relies on static datasets, which cannot meet the real-time personalization needs of users.
Figure 1: Schematic illustration of CoSteer framework. (a) Task scenario: A user poses a question potentially requiring access to local personal context (e.g., user profile, interaction history). (b) Limitations of small locally-deployed language models: Direct inference with constrained model capability leads to suboptimal generation quality. (c) Challenges of cloud-based LLMs: Despite strong generalization, once LLMs are constrained from accessing local personal context, they result in misaligned or contextually disconnected outputs. (d) CoSteer: Optimizes LLM predictions through local delta steering, balancing the LLM’s broad knowledge with user-specific information.
Inference-time Adaptation.
@@
In order to bypass the prohibitive retraining costs, inference-time adaptation has emerged as a lightweight alternative that modulates model behavior dynamically during decoding process. Unlike the conventional practice of augmenting the input prompt with retrieved user profiles or historical interactions(pag; cuecot), methods like Linear Alignment (gao2024linearalignmentclosedformsolution) and Contrastive Decoding (li2023contrastivedecodingopenendedtext) pioneer logit steering mechanisms that dynamically adjust token distributions during generation. By computing differential signals on-device, these approaches eliminate dependency on retraining while preserving privacy through localized computation. CoS (he2025contextsteeringcontrollablepersonalization) follows this approach by using an adaptable parameter $\lambda$ to scale the impact of personal information, thereby achieving controllable personalization. Amulet (zhang2025amulet) uses online learning algorithms to iteratively optimize the logits distribution, precisely aligning with the user’s preferences. However, this series of methods requires explicitly transmitting user information to the LLM, which risking privacy leakage (YAN2025100300) and hinders the practical deployment of these methods.
Collaborative Generation.
@@
The computational constraints of edge devices have driven innovative paradigms in collaborative generation between cloud-based LLMs and local SLMs. Existing collaborative generation approaches predominantly focus on two technical routes: (1) assistive alignment, where local SLMs are trained to augment the LLM’s user preference alignment during inference, and (2) reward-guided decoding, where lightweight reward models provide preference signals through reward-guided decoding. For instance, Aligner (aligner) leverages natural language feedback generated by the SLM to directly inject user-specific linguistic patterns into the LLM’s decoding process, while Expo (expo) achieves preference alignment through linear interpolation of layer-wise weights between LLMs and SLMs. Recent advancements further demonstrate that training specialized small reward models can effectively steer LLM outputs toward personalized objectives (reward1; reward2). Among these, Proxy-tuning (liu2024tuninglanguagemodelsproxy) and Cogensis (cogenesis) are closest to our design. Proxy-tuning modifies the LLM’s logits distribution by contrasting pre-trained and fine-tuned SLM outputs, whereas Cogensis employs a learned fusion network to combine logits from both models. In contrast to these existing methods that rely on parameter optimization, our proposed CoSteer framework establishes a completely tuning-free collaboration mechanism, thereby thoroughly eliminating the need for additional training.
@@ 3.1.1 Task formulation
We consider personalized text generation under a strict edge–cloud deployment: a cloud-hosted LLM only processes the raw user query $p_{base}$, while a locally deployed SLM has access to privacy-sensitive personal context $p_{pers}$ (e.g., user profiles and interaction histories) that must remain on-device. Our goal is to improve generation quality by combining the LLM’s general capability with localized personalization, without transmitting $p_{pers}$ to the cloud.
To enable test-time adaptation under the privacy boundary, we model decoding as a sequence of per-token online optimization problems. At each decoding step with current prefix $s$, the model chooses the next token $a\in\mathcal{A}$ according to a policy $\pi(\cdot\mid p,s)$, where the policy is parameterized by the LLM logit distribution. The per-token objective is:
$\pi^{*}=\underset{\pi\in\Pi}{\arg\max}\ \mathbb{E}_{a\sim\pi\left(\cdot\mid p,s\right)}\big[r\left(a\mid p,s\right)\big],$ (1)
where $p$ denotes a generic prompt, $s$ is the already generated prefix, and $r$ is an (unknown) reward function reflecting the current user’s personal context and preferences.
@@ 3.1.2 Online learning
Personalization signals are individual-specific and may change over time, making purely offline training on static data insufficient for test-time adaptation. Online learning provides a natural abstraction: at each step, the algorithm updates the policy using newly observed utility signals, aiming to minimize regret against the best fixed policy in hindsight.
Inspired by Amulet (zhang2025amulet), we adopt Follow-The-Regularized-Leader (FTRL) (pmlr-v15-mcmahan11b) to refine the token policy during decoding. In its generic form, FTRL updates the policy at iteration $t$ by optimizing the cumulative utilities with an explicit regularizer:
$\pi_{t}=\arg\max_{\pi\in\Pi}\left[\sum_{i=0}^{t-1}\mathcal{U}_{i}(\pi)-\frac{1}{\eta}\Omega(\pi)\right],$ (2)
where $\mathcal{U}_{i}$ is the utility at update iteration $i$, $\Omega(\cdot)$ is a regularizer, and $\eta$ controls the update strength. In our setting, the true reward $r$ in Eq. 1 is not directly observable at inference time, so $\mathcal{U}_{i}$ must be constructed from an efficient proxy signal derived from available information. Throughout the paper, $t$ denotes the current inner FTRL update iteration within a single decoding step, while $i$ indexes the past inner iterations in the cumulative sum; the decoding token position is omitted for simplicity.
@@ 3.2 Decoding-time personalization via local delta steering
To instantiate the online optimization at decoding time, we need a utility function that can be evaluated under a strict edge–cloud privacy boundary: the cloud-hosted LLM only conditions on the raw query $p_{base}$, while the privacy-sensitive personal context $p_{pers}$ must remain on-device. We therefore design a logit-level steering signal that is computed locally and can be used to refine the token distribution without exposing $p_{pers}$.
A series of recent works, including Contrastive Decoding (li2023contrastivedecodingopenendedtext) and Linear Alignment (gao2024linearalignmentclosedformsolution), show that contrasting logits with and without a certain context provides an effective direction for improving generation (he2025contextsteeringcontrollablepersonalization; zhang2025amulet). From our online learning view, such a contrast can be written as the following utility:
$u_{t}(a)=\alpha\left(\log\pi_{t}(a)-\log\pi_{\text{base }}(a)\right).$ (3)
However, directly computing this contrast on the cloud LLM would require transmitting $p_{pers}$ to the cloud, which violates our privacy constraint. Our key observation is that the cloud LLM and the on-device SLM share the same vocabulary and decode under the same prefix $s$, so a context-induced delta computed by the SLM can be used as a compatible steering signal for the LLM logit space. We thus define a dual-contrastive utility as follows:
@@
$\begin{split}u_{t}(a)&=\underbrace{\alpha(\log\pi_{t}(a)-\log\pi_{base}(a))}_{\text{LLM Policy Contrast}}\\ &\quad+\underbrace{\beta(\log\pi^{*}_{pers}(a)-\log\pi^{*}_{base}(a))}_{\text{SLM Delta Steering}}.\end{split}$ (4)
Here, $p_{base}$ is the user query and $p_{pers}$ is the private personal context. We formalize three policies: (1) the target policy $\pi_{t}(a)$ to be optimized; (2) the LLM base policy $\pi_{base}(a)=P_{\text{LLM}}(a\mid p_{base},s)$; and (3) the SLM reference policies $\pi^{*}_{base}(a)=P_{\text{SLM}}(a\mid p_{base},s)$ and $\pi^{*}_{pers}(a)=P_{\text{SLM}}(a\mid p_{base},p_{pers},s)$. The first term encourages controlled deviation from the LLM base behavior, while the second term injects personalization through the on-device delta $\log\pi^{*}_{pers}(a)-\log\pi^{*}_{base}(a)$.
To retain the LLM’s general capability and prevent overly aggressive steering, we add a KL constraint to the objective:
$\mathcal{U}_{t}(\pi)=u_{t}(\pi)-\lambda D_{\mathrm{KL}}\left(\pi\|\pi_{0}\right),$ (5)
where we set the reference policy as $\pi_{0}=\pi_{base}$. Using KL divergence also naturally matches the regularization form required by FTRL, resulting in a proximal-style update:
$\pi_{t}=\underset{\pi\in\Pi}{\arg\max}\left[\sum_{i=0}^{t-1}\mathcal{U}_{i}(\pi)-\frac{1}{\eta}D_{\mathrm{KL}}\left(\pi\|\pi_{t-1}\right)\right].$ (6)
Although Eq. 6 suggests an iterative refinement, performing expensive inner-loop optimization during inference hinders interactive personalization. We therefore derive a closed-form solution, yielding an efficient logit-level fusion rule:
$\begin{split}\pi_{t}(a)\propto\exp\Bigg(&\frac{1}{t\lambda+\frac{1}{\eta}}\bigg(\sum_{i=0}^{t-1}u_{i}(a)\\ &+t\lambda\log\pi_{0}(a)+\frac{1}{\eta}\log\pi_{t-1}(a)\bigg)\Bigg).\end{split}$ (7)
The derivation is provided in Appendix A.2. Equation 7 can be implemented efficiently as element-wise operations in log-probability space followed by a single normalization, making it suitable for on-device execution.
We refer to the resulting edge–cloud decoding procedure—where the device uses the SLM delta and Eq. 7 to select each token while the cloud LLM is fixed—as CoSteer. The concrete inference pipeline is described in Section $3.$
3.3 Edge–cloud inference pipeline
With the closed-form fusion rule in Eq. 7, we now describe how it is instantiated in our target edge–cloud setting. The cloud-hosted LLM only conditions on the raw query $p_{base}$, while the privacy-sensitive personal context $p_{pers}$ is stored and used exclusively on the user device. At each decoding step with current prefix $s$, the cloud returns the LLM base logits (equivalently, $\pi_{base}(\cdot)=P_{LLM}(\cdot\mid p_{base},s)$), and the device performs the personalization and token selection locally.Algorithm 1 summarizes the overall procedure.
Concretely, the device runs the local SLM twice to obtain a pair of reference distributions: $\pi^{*}_{base}(\cdot)=P_{SLM}(\cdot\mid p_{base},s)$ using the query only, and $\pi^{*}_{pers}(\cdot)=P_{SLM}(\cdot\mid p_{base},p_{pers},s)$ using the full personal context. Their log-probability difference $\log\pi^{*}_{pers}-\log\pi^{*}_{base}$ forms the local delta signal in Eq. 4. Once the device receives the cloud logits, it combines the LLM contrast term and the SLM delta term to construct $u_{t}(\cdot)$, and then updates the target policy $\pi_{t}(\cdot)$ using the closed-form update in Eq. 7. The next token is sampled (or greedily selected) from $\pi_{t}$ and sent back to the cloud to continue decoding.
This design enforces a strict privacy boundary: the cloud never observes the sensitive context $p_{pers}$, the raw SLM logits, or their differences; it only receives the final discrete token at each step. Since the fusion happens fully on-device, CoSteer also admits optional local safeguards (e.g., lightweight PII filters) as a drop-in post-processing step before transmission. Overall, CoSteer realizes decoding-time personalization by combining a local delta signal with an efficient closed-form fusion rule, while keeping both the cloud LLM and the privacy boundary unchanged.
Algorithm 1 CoSteer Framework
[…]